# An Axiomatic Risk-Reward Framework for Sustainable Investing

### v3

@@ 5 Empirical analysis
For each of the stocks, we compute a set of daily ESG risk measures and ESG reward–risk ratios, considering a range of values of $\lambda$. We then use the measures to rank the stocks in the dataset. In particular, we compute $\text{ESG-}\text{AVaR}_{\lambda,\tau}$ (with $\tau$ equal to 0.95 and 0.99), the ESG standard deviation ($\text{ESG-}\sigma_{\lambda}$), the ESG-mean ($\text{ESG-}\mu_{\lambda}$), the ESG Rachev ratio (with $\tau$ equal to 0.95 and 0.99), the ESG STAR ratio (with $\tau$ equal to 0.95 and 0.99), and the ESG Sharpe ratio. The risk measures are computed using a historical simulation approach, employing all the observations in the period 2020-2023. Figure 3 displays the ordering of the assets: each company is color-coded according to the industrial sector in which it operates, and the companies in the upper part of the plot are those with the highest value for the indicator. On the horizontal axis, we vary the investor’s $\lambda$ from zero to one. Tables 4 and 5 in Appendix D report the names of the companies in the top and bottom five positions for a selected sample of indicators with $\lambda\in\{0,0.25,0.5,0.75,1\}$.222222 For brevity we report only the measures for the 95th quantile. The complete rankings and the results for the 99th quantile are available upon request.
We make the following observations:232323 We emphasize that the range of $\lambda$ that we considered is quite extreme, and realistically an investor would choose a small value of $\lambda$ to generate only minor variations compared to traditional market portfolios focused only on monetary considerations.
  • ESG-AVaR tends to give similar rankings for similar levels of $\lambda$ up to 0.5; then, the ranking changes significantly and converges to a significantly different ranking for $\lambda=1$ compared to $\lambda=0$. We can explain this behavior by taking into consideration that the ESG scores are characterized by a smaller variability over time; hence the tails of the distributions of returns is the most relevant determinant of the ESG-AVaR (and thus the ranking) even for relatively large values of $\lambda$. For values of $\lambda$ close to 1, the ESG component becomes dominant in the determination of the ranking. The results are similar for $\tau=0.95$ and $\tau=0.99$; the main difference is that for the highest value of $\tau$, the monetary component remains dominant for higher values of $\lambda$.
@@
  • The ranking according to the STAR ratio shows significant changes when $\lambda$ changes. This can be attributed in large part to the fact that the numerator of the ratio ($\text{ESG-}\mu_{\lambda}$) changes significantly with $\lambda$.
@@
Overall, the example shows the power and flexibility of the proposed framework, which can be used to develop empirical analyses both to answer theoretical questions and to implement viable investment strategies suitable for ESG-oriented investors. Further empirical research is required for more extensive results, in particular for the modelization of the joint evolution of returns and ESG using stochastic process.
@@ 6 Conclusions
Individuals and institutions are increasingly aware of the non-monetary impact of their investments, and many are willing to structure their portfolios considering not only on monetary risk and gains, but also the environmental, social, and governance implications. This shift in investors’ goals challenges the traditional financial literature, requiring the development of new analytical tools to describe the behavior of ESG-oriented investors. Our work contributes to the literature by introducing ESG-coherent risk measures, a framework for the measurement of risk for investors with both monetary and ESG goals that provides an axiomatic definition in which risk is measured as a function of a bivariate random variable. The investor in our approach is still rational but follows a multi-dimensional evaluation that considers not only the monetary part but also sustainability. This framework extends the traditional coherent risk measures approach of Artzner et al., (1999).
@@ Appendix A ESG-coherent reward measures
Similarly to how we defined ESG-coherent risk measures, we define ESG-coherent reward measures. As discussed by Rachev et al., (2008), in the univariate case, a reward measure can be defined as any functional defined on the space of the random variables of interest, that is, it should be isotonic with market preferences. Still, it is useful to formalize axiomatically the characteristics of reward measures. We now introduce axioms for multivariate reward measures inspired by the ones for the univariate case discussed in Rachev et al., (2008). Considering a probability space $(\Omega,\mathcal{F},P)$ and a set of bivariate random variables $\mathcal{X}_{2}$, we can define an ESG-adjusted reward measure $\boldsymbol{\theta}_{\lambda}(X_{T}):\mathcal{X}_{2}\rightarrow\mathbb{R}\cup\{\pm\infty\}$, where $\lambda\in[0,1]$ is then ESG-coherent if the following axioms are satisfied:
@@
The translation invariance property follows from SUP-M+ and LH-M+
@@ Appendix B Proof of dual representation of ESG-coherent risk measures
Appendix C Proofs of dual representation for ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}$ and ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}$
@@
Dual representation for ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}$
Here, we prove that ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}(X_{T})$ has the following dual representation:
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}(X_{T})=\sup_{[\zeta_{1}\;\zeta_{2}]^{\prime}\in\mathcal{A}_{{ESG}\text{-}\text{AVaR}_{\lambda,\tau}}}\mathbb{E}[Z_{1}\zeta_{1}+Z_{2}\zeta_{2}],\quad\text{with}$
$\mathcal{A}_{{ESG}\text{-}\text{AVaR}_{\lambda,\tau}}=\left\{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{L}_{\infty}(\Omega,\mathcal{F},P;\mathbb{R}^{2}):\;[\zeta_{1},\;\zeta_{2}]^{\prime}=\xi[1-\lambda,\;\lambda]^{\prime};\;0\leq\xi\leq\frac{1}{1-\tau}\;\text{a.s. }\mathbb{E}[\xi]=1\right\}.$
@@
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}(X_{T})$ $=\min_{\beta\in\mathbb{R}}\Big{\{}\frac{1}{1-\tau}\mathbb{E}\big{[}\big{(}\beta-((1-\lambda)r+\lambda{ESG})\big{)}_{+}\big{]}-\beta\Big{\}},$
@@
Using the rules of subdifferential calculus and Strassen’s theorem (Strassen,, 1965), we get
@@
where $\boldsymbol{\rho}(Z)={ESG}\text{-}\text{AVaR}_{\lambda,\tau}(X_{T})$. The exchange of the $\min$” and $\max$” operations is possible, because the function in braces is bilinear in $(\beta,\zeta)$, and the set $\tilde{\mathcal{A}}$ is compact. The inner minimization with respect to $\beta$ yields -$\infty$, unless $\mathbb{E}[\xi]=1-\tau$. Consequently,
@@
Dual representation for ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}$
We now prove that the dual representation of ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}(X_{T})$ is defined as follows:
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}(X_{T})=\sup_{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{A}_{{ESG}\text{-}\text{AVaR}_{\lambda,\tau}}}\mathbb{E}[Z_{1}\zeta_{1}+Z_{2}\zeta_{2}],$ (39)
@@
$\mathcal{A}_{{ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}}=\left\{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{L}_{\infty}(\Omega,\mathcal{F},P;\mathbb{R}^{2}):\mathbb{E}[\zeta_{1}]=1-\lambda;\mathbb{E}[\zeta_{2}]=\lambda;\zeta_{1},\zeta_{2}\geq 0;\zeta_{1}\leq\frac{1-\lambda}{1-\tau};\zeta_{2}\leq\frac{\lambda}{1-\tau}\right\}.$ (40)
We use the known representation of Average Value at Risk for scalar random variables (cf. Shapiro et al.,, 2021, Example 6.19 eq. 6.76). Denote the dual set in that representation by $\mathcal{A}^{\prime}$, i.e.,
@@
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}(X)$ $=(1-\lambda)\sup_{\xi\in\mathcal{A}^{\prime}}\mathbb{E}[\xi,Z_{1}]+\lambda\sup_{\xi\in\mathcal{A}^{\prime}}\mathbb{E}[\xi,Z_{2}]$
@@
Now, we define the set $\mathcal{A}_{{ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}}=(1-\lambda)\mathcal{A}^{\prime}\times\lambda\mathcal{A}^{\prime}$ and continue the last chain of equations as follows:
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}(X_{T})$ $=\sup_{[\xi_{1},\xi_{2}]^{\prime}\in\mathcal{A}^{\prime}\times\mathcal{A}^{\prime}}\mathbb{E}[(1-\lambda)\xi,Z_{1}]+\mathbb{E}[\lambda\xi,Z_{2}]=\sup_{\zeta\in\mathcal{A}_{{ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}}}\langle\zeta,Z\rangle.$
@@
Appendix D Ranking of companies based on ESG risk measures, reward measures, and RRRs.
@@ 28 Johnson & Johnson Johnson & Johnson Johnson & Johnson Johnson & Johnson Coca-Cola Company
Table 4: Top and bottom ranking according to several ESG risk and reward measures ($\text{ESG-}\text{AVaR}_{\lambda,0.95\%}$, ESG-standard deviation, ESG-mean) for investors with different $\lambda$.
@@ 25 IBM Corporation IBM Corporation 3M Company Johnson & Johnson Coca-Cola Company
Table 5: Top and bottom ranking according to several ESG-RRRs (ESG Sharpe ratio, $\text{ESG-}\text{STARR}_{\lambda,0.95\%}$, $\text{ESG-}\text{RR}_{\lambda,0.95\%}$) for investors with different $\lambda$.

### v4

@@ 5 Empirical analysis
For each of the stocks, we compute a set of daily ESG risk measures and ESG reward–risk ratios, considering a range of values of $\lambda$. We then use the measures to rank the stocks in the dataset. In particular, we compute $\text{ESG-}\text{AVaR}_{\lambda,\tau}$ (with $\tau$ equal to 0.95 and 0.99), the ESG standard deviation ($\text{ESG-}\sigma_{\lambda}$), the ESG-mean ($\text{ESG-}\mu_{\lambda}$), the ESG Rachev ratio (with $\tau$ equal to 0.95 and 0.99), the ESG STAR ratio (with $\tau$ equal to 0.95 and 0.99), and the ESG Sharpe ratio. The risk measures are computed using a historical simulation approach, employing all the observations in the period 2020-2023. The range of $\lambda$ considered is intentionally extreme and serves as a stress test for the framework. In practice, investors would typically choose smaller values of $\lambda$, resulting in portfolios that remain closer to conventional market portfolios. Figure 3 displays the ordering of the assets: each company is color-coded according to the industrial sector in which it operates, and the companies in the upper part of the plot are those with the highest value for the indicator. On the horizontal axis, we vary the investor’s $\lambda$ from zero to one. Tables 4 and 5 in Appendix D report the names of the companies in the top and bottom five positions for a selected sample of indicators with $\lambda\in\{0,0.25,0.5,0.75,1\}$. [^20] For brevity we report only the measures for the 95th quantile. The complete rankings and the results for the 99th quantile are available upon request.
We make the following observations:
  • ESG-AVaR tends to give similar rankings for levels of $\lambda$ up to 0.5; then, the ranking changes significantly and converges to a significantly different ranking for $\lambda=1$ compared to $\lambda=0$. We can explain this behavior by taking into consideration that the ESG scores are characterized by a smaller variability over time; hence the tails of the distributions of returns is the most relevant determinant of the ESG-AVaR (and thus the ranking) even for relatively large values of $\lambda$. For irrealistically large values of $\lambda$ close to 1, the ESG component becomes dominant in the determination of the ranking. The results are similar for $\tau=0.95$ and $\tau=0.99$; the main difference is that for the highest value of $\tau$, the monetary component remains dominant for higher values of $\lambda$.
@@
  • The ranking according to the STAR ratio shows significant changes when $\lambda$ changes. This can be attributed in large part to the fact that the numerator of the ratio ($\text{ESG-}\mu_{\lambda}$) by construction changes significantly with $\lambda$.
@@
Overall, the example shows the structural flexibility and interpretability of the proposed framework, which can be used to develop empirical analyses both to answer theoretical questions and to implement viable investment strategies suitable for ESG-oriented investors. Further empirical research is required for more extensive results, in particular for the modelization of the joint evolution of returns and ESG using stochastic process and the extension to more assets and geographical areas. Another natural extension is a benchmarking of the results against more ESG data providers, especially considering that the ESG scores from different rating agencies are often inconsistent (Billio et al., 2021).
6 Policy & Practical Implications
Modeling investment outcomes as a bivariate random vector $X_{T}=[r_{T},\;{ESG}_{T}]^{\prime}$, with $r_{T}$ denoting financial payoffs and ${ESG}_{T}$ the sustainability dimension, extends the standard univariate risk and reward measure framework in a structurally meaningful way. ESG-coherent risk measures embed sustainability considerations directly into the risk functional through a preference parameter $\lambda$, allowing ESG to enter portfolio choice in the same formal manner as other preference-driven dimensions. This provides a unified setting in which financial and ESG components are evaluated jointly, rather than sequentially or through external constraints.
For optimal asset allocation, this formulation yields well-defined optimization problems in which portfolio choice depends on the joint distribution of $X_{t}$, with only minimal modifications relative to standard univariate setups. ESG preferences are reflected through $\lambda$, inducing smooth adjustments in optimal portfolios reflecting the subjective trade-off between the monetary and the sustainability dimensions. Unlike exclusion or screening approaches, this mechanism preserves diversification properties and allows for a systematic analysis of sensitivity with respect to ESG preferences, dependence between $r_{T}$ and ${ESG}_{T}$, and joint tail behavior. From a theoretical perspective, it opens the way to studying ESG-aware allocation using familiar tools from coherent risk theory and convex optimization.
The framework is deliberately focused on ESG-oriented investing, in the sense that ESG enters the model as an argument of investor preferences rather than as a predictive signal for financial returns. This does not preclude the use of ESG information for improving financial performance, for example through forecasting, risk premia estimation, or alpha generation. Rather, the advantage of the proposed approach is to keep these two roles of ESG analytically distinct: ESG can be treated either as a component of the payoff vector $X_{T}$ reflecting non-financial objectives, or as an informational variable influencing the distribution of $r_{T}$. By separating preference-driven valuation from return-predictive use, the framework avoids conflating ethical objectives with performance claims and clarifies the economic interpretation of ESG integration, while remaining flexible enough to accommodate their joint use within advanced asset allocation strategies designed to address multiple, potentially heterogeneous, investment objectives.
The stochastic treatment of ${ESG}_{T}$ also has direct implications for ESG data provision. Meaningful ESG-coherent risk measurement requires time-series ESG data at frequencies compatible with financial modeling, creating incentives for data providers to move beyond static or purely cross-sectional scores. Higher-frequency and methodologically consistent ESG measures directly improve estimation of the joint law of $X_{T}=[r_{T},\;{ESG}_{T}]^{\prime}$, its dependence structure, and joint tail risk, rather than merely refining descriptive asset rankings.
In terms of transparency and reporting, ESG-coherent risk measures make the aggregation of financial and ESG outcomes explicit. The role of preferences is isolated in the parameter $\lambda$, which facilitates comparability across portfolios and products by separating differences due to investor objectives from differences due to underlying risk exposures. For institutional investors, this supports internal coherence between mandates, portfolio construction, and performance evaluation. For retail investors, it clarifies the trade-offs embedded in ESG-oriented products, reducing reliance on implicit or opaque screening rules.
Finally, the framework makes explicit several structural limitations that are often implicit in ESG integration. The choice of $\lambda$ is normative and cannot be identified from data alone, and the quality of ESG-coherent risk estimates is constrained by data availability and by the modeling of dependence between $r_{T}$ and ${ESG}_{T}$. Making these elements explicit shifts attention from the proliferation of ESG scores to the underlying statistical and preference-based structure of risk measurement for ESG-oriented investors, where further theoretical and empirical progress is most likely to be achieved.
@@ 7 Conclusions
Individuals and institutions are increasingly aware of the non-monetary impact of their investments, and many are willing to structure their portfolios considering not only monetary risk and gains, but also the environmental, social, and governance implications. These ESG-oriented investors, who are rational under a richer set of objectives, challenge the traditional financial literature and require the development of new analytical tools to describe their behavior. Our work contributes to the literature by introducing ESG-coherent risk measures, a framework for the measurement of risk for investors with both monetary and ESG goals that provides an axiomatic definition in which risk is measured as a function of a bivariate random variable. The investor in our approach is still rational but follows a multi-dimensional evaluation that considers not only the monetary part but also sustainability. This framework extends the traditional coherent risk measures approach of Artzner et al. (1999).
@@ Appendix A ESG-coherent reward measures
Similarly to how we defined ESG-coherent risk measures, we define ESG-coherent reward measures. As discussed by Rachev et al. (2008), in the univariate case, a reward measure can be defined as any functional defined on the space of the random variables of interest, that is, it should be isotonic with market preferences. Still, it is useful to formalize axiomatically the characteristics of reward measures. We now introduce axioms for multivariate reward measures inspired by the ones for the univariate case discussed in Rachev et al. (2008). Considering a probability space $(\Omega,\mathcal{F},P)$ and a set of bivariate random variables $\mathcal{X}_{2}$, we can define an ESG-adjusted reward measure $\boldsymbol{\theta}_{\lambda}(X_{T}):\mathcal{X}_{2}\rightarrow\mathbb{R}\cup\{\pm\infty\}$, where $\lambda\in[0,1]$ is then ESG-coherent if the following axioms are satisfied:
@@
The translation invariance property follows directly from SUP-M+ and LH-M+, analogously to Proposition 1.
@@ Appendix B Proof of dual representation of ESG-coherent risk measures
Appendix C Proofs of dual representation for $\text{ESG-}\text{AVaR}_{\lambda,\tau}$ and $\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}$
@@
Dual representation for $\text{ESG-}\text{AVaR}_{\lambda,\tau}$
Here, we prove that $\text{ESG-}\text{AVaR}_{\lambda,\tau}(X_{T})$ has the following dual representation:
$\text{ESG-}\text{AVaR}_{\lambda,\tau}(X_{T})=\sup_{[\zeta_{1}\;\zeta_{2}]^{\prime}\in\mathcal{A}_{\text{ESG-}\text{AVaR}_{\lambda,\tau}}}\mathbb{E}[Z_{1}\zeta_{1}+Z_{2}\zeta_{2}],\quad\text{with}$
$\mathcal{A}_{\text{ESG-}\text{AVaR}_{\lambda,\tau}}=\left\{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{L}_{\infty}(\Omega,\mathcal{F},P;\mathbb{R}^{2}):\;[\zeta_{1},\;\zeta_{2}]^{\prime}=\xi[1-\lambda,\;\lambda]^{\prime};\;0\leq\xi\leq\frac{1}{1-\tau}\;\text{a.s. }\mathbb{E}[\xi]=1\right\}.$
@@
$\text{ESG-}\text{AVaR}_{\lambda,\tau}(X_{T})$ $=\min_{\beta\in\mathbb{R}}\Big\{\frac{1}{1-\tau}\mathbb{E}\big[\big(\beta-((1-\lambda)r+\lambda{ESG})\big)_{+}\big]-\beta\Big\},$
@@
Using the rules of subdifferential calculus and Strassen’s theorem (Strassen, 1965), we get
@@
where $\boldsymbol{\rho}(Z)=\text{ESG-}\text{AVaR}_{\lambda,\tau}(X_{T})$. The exchange of the $\min$” and $\max$” operations is possible, because the function in braces is bilinear in $(\beta,\zeta)$, and the set $\tilde{\mathcal{A}}$ is compact. The inner minimization with respect to $\beta$ yields -$\infty$, unless $\mathbb{E}[\xi]=1-\tau$. Consequently,
@@
Dual representation for $\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}$
We now prove that the dual representation of $\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}(X_{T})$ is defined as follows:
$\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}(X_{T})=\sup_{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{A}_{\text{ESG-}\text{AVaR}_{\lambda,\tau}}}\mathbb{E}[Z_{1}\zeta_{1}+Z_{2}\zeta_{2}],$ (39)
@@
$\mathcal{A}_{\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}}=\left\{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{L}_{\infty}(\Omega,\mathcal{F},P;\mathbb{R}^{2}):\mathbb{E}[\zeta_{1}]=1-\lambda;\mathbb{E}[\zeta_{2}]=\lambda;\zeta_{1},\zeta_{2}\geq 0;\zeta_{1}\leq\frac{1-\lambda}{1-\tau};\zeta_{2}\leq\frac{\lambda}{1-\tau}\right\}.$ (40)
We use the known representation of Average Value at Risk for scalar random variables (cf. Shapiro et al., 2021, Example 6.19 eq. 6.76). Denote the dual set in that representation by $\mathcal{A}^{\prime}$, i.e.,
@@
$\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}(X)$ $=(1-\lambda)\sup_{\xi\in\mathcal{A}^{\prime}}\mathbb{E}[\xi,Z_{1}]+\lambda\sup_{\xi\in\mathcal{A}^{\prime}}\mathbb{E}[\xi,Z_{2}]$
@@
Now, we define the set $\mathcal{A}_{\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}}=(1-\lambda)\mathcal{A}^{\prime}\times\lambda\mathcal{A}^{\prime}$ and continue the last chain of equations as follows:
$\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}(X_{T})$ $=\sup_{[\xi_{1},\xi_{2}]^{\prime}\in\mathcal{A}^{\prime}\times\mathcal{A}^{\prime}}\mathbb{E}[(1-\lambda)\xi,Z_{1}]+\mathbb{E}[\lambda\xi,Z_{2}]=\sup_{\zeta\in\mathcal{A}_{\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}}}\langle\zeta,Z\rangle.$
@@
Appendix D Company rankings based on ESG risk, reward, and RRR measures
Table 4: Top and bottom ranking according to several ESG risk and reward measures ($\text{ESG-}\text{AVaR}_{\lambda,0.95\%}$, ESG-standard deviation, ESG-mean) for selected values of $\lambda$.
@@ 28 Johnson & Johnson Johnson & Johnson Johnson & Johnson Johnson & Johnson Coca-Cola Company
Table 5: Top and bottom ranking according to several ESG-RRRs (ESG Sharpe ratio, $\text{ESG-}\text{STARR}_{\lambda,0.95\%}$, $\text{ESG-}\text{RR}_{\lambda,0.95\%}$) for selected values of $\lambda$.